OpenStudy (anonymous):

willing to give medals and become if somebody could help me answer this question

OpenStudy (anonymous):

Solve 6 over x minus 6 equals the quantity of x over x minus 6, minus six halfs for x and determine if the solutions is extraneous or not. for x and determine if the solution is extraneous or not.

OpenStudy (anonymous):

here are the answer choices x = -6, extraneous x = -6, non-extraneous x = 6, non-extraneous x = 6, extraneous

OpenStudy (nikato):

Ur first step would be making all the fractions have a common denominator To this multiply the first and second each with 2/2 Multiply the last one by (x-6)/(x-6)

OpenStudy (anonymous):

ok well the first fraction would be 2x/2x-12 right @nikato

OpenStudy (nikato):

Yes, that's one of them

OpenStudy (anonymous):

and would the second fraction be 6/x-6

OpenStudy (anonymous):

right

OpenStudy (nikato):

No. This is getting confusing.. 6 2. x. 2. 6. (x-6) ---- * ----- = ------ *. ------ - ---- * ----- (x-6). 2. (x-6). 2. 2. (x-6) Does this make sense? If so, can you simplify the numerator, because the denominator is all the same

OpenStudy (anonymous):

yea kinda

OpenStudy (nikato):

So basically, I mutinied the first and second by 2/2 and the lat by (x-6)/(x-6)

OpenStudy (nikato):

*multiplied. And now all 3 fractions share a common denominator, 2(x-6)

OpenStudy (anonymous):

ok

OpenStudy (nikato):

Since we know our denominator is 2(x-6), what can't x be? Remember the denominator can't =0

OpenStudy (anonymous):

-6?

OpenStudy (nikato):

No, 6. 2(x-6)=0 ---- -- 2. 2 x-6=0 +6. +6 ----------- x=6

OpenStudy (anonymous):

so its extraneous

OpenStudy (nikato):

Ok, we can just ignore the denominator now and work on the numerator. 6(2)= x(2) - 6(x-6) Simplify and solve for x

OpenStudy (anonymous):

ok i solved for it but was i right when i said it was extraneous @nikato

OpenStudy (nikato):

Oh ok. Yea. 6 is extraneous